OpenStudy (anonymous):

(x2 + 4x – 3)(2x2 + x + 6)??

OpenStudy (anonymous):

We'll use FOIL method (x^2 + 4x - 3)(2x^2 + x + 6) = x^2*2x^2 + x^2*x + x^2*6 + 4x*2x^2 + 4x*x + 4x*6 - 3*2x^2 - 3x - 18 (x^2 + 4x - 3)(2x^2 + x + 6) =x^4 + x^3 + 6x^2 + 8x^3 + 4x^2 + 24x - 6x^2 - 3x - 18 (x^2 + 4x - 3)(2x^2 + x + 6) =x^4 +9x^3 + 4x^2+21x - 18